Purple Rain Company is developing a new technology that has the potential of forever changing the way that vegetables are grown. The initial costs associated with its manufacturing facility plant and other associated costs will be $3.2 million (at the present time).

The project is expected to generate only one cash flow of $8.3 million. Given the complexity of the project, the cash inflow will not be received until 10 years from now.

Calculate the internal rate of return on the investment.
Answer #1

IRR = [Cash Inflow in 10 years / Cash Outflow]1/n - 1

= [$8.3 million / $3.2 million]1/10 - 1

= [2.59375]0.1 - 1 = 1.10 - 1 = 0.10, or 10%